OpenStudy (anonymous):

what is the solution to this system of linear equations? y=2x-4 y=x+3 A. NO SOLUTION B. (3,2) C.(5,6) D. (7,10)

OpenStudy (matt101):

x + 3 = 2x - 4 -x = -7 x = 7 y = x + 3 y = 7 + 3 y = 10 Answer is D